# Three-in-One: Fast and Accurate Transducer for Hybrid-Autoregressive ASR

### v1

@@ Abstract
We present Hybrid-Autoregressive INference TrANsducers (HAINAN), a novel architecture for speech recognition that extends the Token-and-Duration Transducer (TDT) model. Trained with randomly masked predictor network outputs, HAINAN supports both autoregressive inference with all network components and non-autoregressive inference without the predictor. Additionally, we propose a novel semi-autoregressive inference paradigm that first generates an initial hypothesis using non-autoregressive inference, followed by refinement steps where each token prediction is regenerated using parallelized autoregression on the initial hypothesis. Experiments on multiple datasets across different languages demonstrate that HAINAN achieves efficiency parity with CTC in non-autoregressive mode and with TDT in autoregressive mode. In terms of accuracy, autoregressive HAINAN outperforms TDT and RNN-T, while non-autoregressive HAINAN significantly outperforms CTC. Semi-autoregressive inference further enhances the model’s accuracy with minimal computational overhead, and even outperforms TDT results in some cases. These results highlight HAINAN’s flexibility in balancing accuracy and speed, positioning it as a strong candidate for real-world speech recognition applications.
@@ 1 Introduction
End-to-end neural automatic speech recognition (ASR) has seen significant advancements in recent years, namely due to the development of three architecture paradigms: Connectionist Temporal Classification (CTC) (Graves et al., 2006), Recurrent Neural Network Transducers (RNN-T) (Graves, 2012), and Attention-based Encoder and Decoder Models (Chorowski et al., 2015; Chan et al., 2016). These models have gained widespread adoption, supported by open-source projects such as ESPNet (Watanabe et al., 2018), SpeechBrain (Ravanelli et al., 2021), and NeMo (Kuchaiev et al., 2019), etc.
CTC and RNN-T models share a frame-synchronous design, enabling streaming processing of speech input. However, they differ in their inference approaches: CTC models employ non-autoregressive (NAR) inference based on a conditional-independence assumption, while RNN-T models require autoregressive (AR) inference due to their dependence on the history context of partial hypotheses. This fundamental difference creates a trade-off between accuracy and efficiency, with CTC models offering greater speed but typically lower accuracy than Transducers.
Our work bridges this gap with Hybrid-Autoregressive INference TrANsducer (HAINAN)111 Despite similar names, HAINAN is unrelated to Hybrid Autoregressive Transducer (HAT) (Variani et al., 2020).: a novel implementation of RNN-T capable of non-autoregressive, autoregressive and, semi-autoregressive inference. Our contributions are as follows:
@@
  • We present a novel semi-autoregressive (SAR) inference paradigm that first uses NAR to generate initial hypotheses, and then uses autoregression to iteratively refine the hypotheses in parallel.
  • We present a means to model the output of the NAR HAINAN model containing the token and duration predictions as a directed acyclic graph (DAG), and propose a simple Viterbi-based decoding method for inference.
@@
  • AR HAINAN achieves better WER accuracy than RNN-T and TDT models. This is achieved with inference speeds similar to TDT-based decoding, and significantly faster than RNN-T inference.
  • NAR HAINAN significantly outperform CTC accuracy, while maintaining similar efficiency.
  • SAR consistently brings accuracy gains over NAR, approaching performance of AR inference with minimal computational overhead.
@@
We open-source our implementation and release trained model checkpoints for public use.
@@ 2 Related Work
Recent research in speech recognition has focused on enhancing WER accuracy while mitigating reductions in inference speed. This has been most notable for self-attention based architectures, with improvements in latency for Transformer models (Yeh et al., 2019) leading to creation of the Conformer (Gulati et al., 2020) and it variants (e.g. Zipformer (Yao et al., 2023), FastConformer (Rekesh et al., 2023)). While all have notable architectural improvements, they can be generalized as attempts to improve model capacity and/or reducing the latency caused by attention-based inference over large audio inputs.
For Transducer models specifically, various optimizations have been explored. He et al. (2019) proposed a caching mechanism for the RNN-T predictor to reduce redundancy in inference computations. Jain et al. (2019) introduced a pruning method for RNN-T beam search to reduce the search space, resulting in faster decoding. Ghodsi et al. (2020) investigated predictors with stateless networks for Transducer models. While this approach reduced the computational complexity of the original LSTM predictor, it came at the cost of slight accuracy degradation. Kang et al. (2023) proposed monotonic constraints on RNN-T training and inference, reducing the number of steps RNN-T inference stays on the same frame, and reported faster inference with those models. Bataev et al. (2024) proposed a novel label-looping decoding algorithm for Transducers, using parallel GPU calls for the majority of the decoding operation and achieved significant inference speedup. Galvez et al. (2024) optimized the decoding algorithm from a hardware perspective and minimized the GPU idle time during decoding, further improving GPU-based inference. Multi-blank Transducers (Xu et al., 2022) and Token-and-Duration Transducers (TDT) (Xu et al., 2023) identified large numbers of blank predictions during RNN-T inference, proposing duration-prediction mechanisms to skip frames during processing. This improvement greatly reduced the number of decoding steps required for inference. While these methods improved Transducer inference speeds, the resulting models remained autoregressive. As such, Transducer inference speeds still significantly lagged behind non-autoregressive frameworks (e.g. CTC).
@@ 3 Background
In this Section, we provide the technical background information for models that influenced HAINAN architecture: Connectionist Temporal Classification (CTC), RNN-Transducer (RNN-T), and Token-and-Duration Transducer (TDT).
@@ 3.1 Connectionist Temporal Classification (CTC)
(a) CTC Architecture
(b) RNN-T Architecture
(c) TDT Architecture
Figure 1: Architecture of CTC, RNN-T and TDT Models.
CTC (Graves et al., 2006) is a sequence-to-sequence prediction model designed to address the challenge of aligning input and output sequences of different lengths. Let $x=(x_{1},x_{2},...,x_{T})$ represent the input sequence, $y=(y_{1},y_{2},...,y_{U})$ the output sequence, and $\mathcal{V}$ the vocabulary, where $v\in\mathcal{V}$ includes all tokens in $y$ and a special “blank” symbol. As illustrated in Figure 1, a CTC model predicts $P(v|x_{t})$ for each time step $t$. The “blank” symbol is crucial for handling the discrepancy between input and output sequence lengths, allowing for many-to-one alignments and representing frames that do not contribute additional information to the output. CTC models employ a set of rules to generate CTC-augmented sequences that are equivalent to the original sequence: (1) When a token is repeated in the original sequence, at least one “blank” token must be inserted between the repeated tokens, and (2) Any token can be repeated an arbitrary number of times.
@@ 3.2 Recurrent Neural Network Transducer (RNN-T)
The RNN-Transducer, proposed by Graves (2012), represents a significant advancement in sequence-to-sequence modeling. As illustrated in Fig. 1, RNN-T comprises three key components: an encoder, a predictor, and a joint network. The encoder extracts high-level features from the audio input, while the predictor processes the text history. The joint network then combines these outputs to generate a probability distribution over the vocabulary.
@@
Despite its superior performance, RNN-T’s reliance on the predictor and text history introduces autoregressive behavior in inference. Consequently, RNN-T inference resists full parallelization, resulting in lower efficiency compared to CTC. This trade-off between performance and efficiency highlights the ongoing challenge in designing optimal sequence-to-sequence models for ASR.
@@ 3.3 Token-and-Duration Transducers (TDT)
Token-and-Duration Transducer (TDT) model is introduced by Xu et al. (2023). Unlike traditional Transducers where only blank predictions increment the encoder index $t$ by exactly one, TDT supports advancing $t$ by multiple steps for both blank and non-blank predictions. As illustrated in Figure 1, a TDT model generates two probability distributions, one for token, and the other for durations, which determines the number of audio frames to skip. By supporting frame skipping, TDT models require significantly fewer decoding steps compared to traditional Transducers, resulting in faster inference. It’s also worth noting that TDT models generally outperform conventional Transducer models in terms of recognition accuracy (Xu et al., 2023).
@@ 4 Hybrid-Autoregressive INference TrANsducer (HAINAN)
We propose Hybrid-Autoregressive INference TrANsducers (HAINAN), an extension of the TDT model designed to support multiple modes of autoregression in inference. HAINAN incorporates stochastic predictor masking: during training, the predictor’s output is randomly masked before being passed to the joint network. We implement this with a masking probability of 0.5, sampled at the [batch, text-index] level. This modification requires a simple one-line change to the existing joint network computation code in TDT model implementation:
         ... # getting f from encoder, g from predictor      f = f.unsqueeze(dim=2)  # encoder output (B, T, 1, H)      g = g.unsqueeze(dim=1)  # decoder output (B, 1, U, H)  +   g = g * torch.gt(torch.rand([B, 1, U, 1]), 0.5)      merged = f + g  # [B, T, U, H]      ... # pass merged to joint network and loss computation
Except for this one-line change, the rest of the model training of HAINAN (computation in other components, and loss computation, etc) remains identical to that of TDT training. As the result of this change, during training, half the time the joint network is learning to predict with the whole model, and the other half the time it is learning to predict with information from the encoder only. Therefore when the model is well-trained, the joint network can generate valid distributions regardless of whether the predictor output is provided, and the model can perform inference in two modes.
@@ 4.1 Semi-autoregressive inference
HAINAN models support three inference modes: autoregressive (AR), non-autoregressive (NAR), and semi-autoregressive (SAR). The AR inference is identical to the TDT’s inference algorithm (see Algorithm 2 of (Xu et al., 2023)), and the NAR inference requires simple modifications of (1) passing an all-zero tensor instead of predictor output to the joint network, and (2) moving the joint network computation out of the decoding loop, allowing parallel processing of all encoder outputs.
Algorithm 1 Semi-Autoregressive Inference of HAINAN Models
1:  Input: acoustic input $x$
2:  enc = encoder(x)
3:  hyp, time-stamps = NAR-Inference(enc)
4:  useful-frames = enc[time-stamps,:]
5:  for i in range(refine-steps) do
6:     shifted-hyp = [BOS] + hyp[:-1]
7:     pred = predictor(shifted-hyp)
8:     token-probs, _ = joint(useful-frames, pred)
9:     hyp = argmax(token-probs, dim=-1) # argmax not including blanks here.
10:  Return: hyp
SAR inference is a novel algorithm combining elements of both AR and NAR approaches, as detailed in Algorithm 1. It consists of two main phases:
@@
    Hypothesis Refinement (lines 5-9) uses the predictor to compute representations of all partial history contexts in the current hypothesis. It then combines the predictor’s output with selected encoder output frames in the joint network to re-compute probability distributions for those frames, in parallel [^2] The computation at line 8 is done fully parallelized, and line 7 may require some sequential computation if the predictor is an LSTM. If the predictor is stateless, this computation can also be fully parallelized as well.. This procedure can be repeated to further improve the results.
The SAR approach leverages the strengths of both AR and NAR methods, enabling high degrees of parallelization while still taking advantage of text history in the predictor computation to improve the accuracy of the model.

### v3

# HAINAN: Fast and Accurate Transducer for Hybrid-Autoregressive ASR

@@ Abstract
We present Hybrid-Autoregressive INference TrANsducers (HAINAN), a novel architecture for speech recognition that extends the Token-and-Duration Transducer (TDT) model. Trained with randomly masked predictor network outputs, HAINAN supports both autoregressive inference with all network components and non-autoregressive inference without the predictor. Additionally, we propose a novel semi-autoregressive inference method that first generates an initial hypothesis using non-autoregressive inference, followed by refinement steps where each token prediction is regenerated using parallelized autoregression on the initial hypothesis. Experiments on multiple datasets across different languages demonstrate that HAINAN achieves efficiency parity with CTC in non-autoregressive mode and with TDT in autoregressive mode. In terms of accuracy, autoregressive HAINAN achieves parity with TDT and RNN-T, while non-autoregressive HAINAN significantly outperforms CTC. Semi-autoregressive inference further enhances the model’s accuracy with minimal computational overhead, and even outperforms TDT results in some cases. These results highlight HAINAN’s flexibility in balancing accuracy and speed, positioning it as a strong candidate for real-world speech recognition applications.
@@ 1 Introduction
End-to-end neural automatic speech recognition (ASR) has seen significant advancements in recent years, namely due to the development of three architecture paradigms: Connectionist Temporal Classification (CTC) (Graves et al., 2006), Recurrent Neural Network Transducers (RNN-T) (Graves, 2012), and Attention-based Encoder and Decoder Models (Chorowski et al., 2015; Chan et al., 2016). These models have gained widespread adoption, supported by open-source projects such as ESPNet (Watanabe et al., 2018), SpeechBrain (Ravanelli et al., 2021), and NeMo (Kuchaiev et al., 2019), etc. Of those models, CTC and RNN-T share a frame-synchronous design, enabling streaming processing of speech input. However, they differ in their inference approaches: CTC models support non-autoregressive (NAR) inference based on a conditional-independence assumption, while RNN-T models require autoregressive (AR) inference due to their dependence on the history context of partial hypotheses. This fundamental difference creates a trade-off between accuracy and efficiency, with CTC models offering greater speed but typically lower accuracy than Transducers.
Our work bridges this gap with Hybrid-Autoregressive INference TrANsducer (HAINAN)111 Despite similar names, our work is unrelated to Hybrid Autoregressive Transducer (Variani et al., 2020).: a novel implementation of RNN-T capable of non-autoregressive, autoregressive and, semi-autoregressive inference. Our contributions are as follows:
@@
  • We present a novel semi-autoregressive (SAR) inference method that first uses NAR to generate initial hypotheses, and then uses autoregression to refine the hypotheses in parallel iteratively.
  • We present a means to model the output of the NAR HAINAN model containing the token and duration predictions as a directed acyclic graph (DAG), and propose a simple Viterbi-based decoding method for inference, with further accuracy gains.
@@
  • AR HAINAN achieves on-par WER accuracy with RNN-T and TDT models, with inference speeds similar to TDT-based decoding, and significantly faster than RNN-T inference.
  • NAR HAINAN outperforms CTC accuracy while maintaining similar efficiency. Our investigation shows that NAR HAINAN can better capture potential ambiguities in the speech by learning more nuanced representations, a capability that frame-by-frame NAR like CTC inherently lacks.
  • SAR consistently brings accuracy gains over NAR, approaching the performance of AR inference with minimal computational overhead.
@@
We will open-source our implementation and release trained model checkpoints for public use.
@@ 2 Related Work
Recent research in speech recognition has focused on enhancing WER accuracy while mitigating reductions in inference speed. This has been most notable for self-attention-based architectures, with improvements in latency for Transformer models (Yeh et al., 2019) leading to creation of the Conformer (Gulati et al., 2020) and it variants (e.g. Zipformer (Yao et al., 2023), FastConformer (Rekesh et al., 2023)). While all have notable architectural improvements, they can be generalized as attempts to improve model capacity and/or reducing the latency caused by attention-based inference over large audio inputs.
For Transducer models specifically, various optimizations have been explored. He et al. (2019) proposed a caching mechanism for the RNN-T predictor to reduce redundancy in inference computations. Jain et al. (2019) introduced a pruning method for RNN-T beam search to reduce the search space, resulting in faster decoding. Ghodsi et al. (2020) investigated predictors with stateless networks for Transducer models. While this approach reduced the computational complexity of the original LSTM predictor, it came at the cost of slight accuracy degradation. Kang et al. (2023) proposed monotonic constraints on RNN-T training and inference, reducing the number of steps RNN-T inference stays on the same frame, and reported faster inference with those models. Bataev et al. (2024) proposed a novel label-looping decoding algorithm for Transducers, using parallel GPU calls for the majority of the decoding operation and achieved significant inference speedup. Galvez et al. (2024) optimized the decoding algorithm from a hardware perspective and minimized the GPU idle time during decoding, further improving GPU-based inference. Multi-blank Transducers (Xu et al., 2022) and Token-and-Duration Transducers (TDT) (Xu et al., 2023b) identified large numbers of blank predictions during RNN-T inference, proposing duration-prediction mechanisms to skip frames during processing. This improvement greatly reduced the number of decoding steps required for inference. While these methods improved Transducer inference speeds, the resulting models remained autoregressive. As such, Transducer inference speeds still significantly lagged behind non-autoregressive frameworks (e.g. CTC).
Our proposed semi-autoregressive method falls under the category of “fast/slow” methods, where the model combines a faster but less accurate module, and a slower but more accurate module to achieve good accuracy/efficiency tradeoffs. Inaguma et al. (2021) use a fast CTC model to generate multiple hypotheses, to be re-ranked by a slower AR model to improve speech translation performance. Mahadeokar et al. (2022) propose a cascade model architecture, where the fast model uses a relatively shallow encoder for quicker results, while the slow model uses additional encoder layers along with a separate joint network to improve results. Arora et al. (2024) proposed a model that divides speech input into blocks, and adopts AR processing among blocks but NAR processing within a block to balance speed with accuracy. Deliberation methods Hu et al. (2020) and its variants Wang et al. (2022) use cascaded encoders and separate decoder, and use the shallow encoder to generate initial hypotheses, which to be fed into a second-pass “Listen-Attend-Spell” (LAS) model with additional encoder and decoders to improve accuracy. Compared with those “fast-slow” works, our work is unique in that our “slow” model only adds marginal parameters on top of the “fast” model, and the “refinement” process requires minimal computation overheads.
@@ 3 Background
In this Section, we provide the technical background for models that influenced HAINAN architecture: CTC, RNN-Transducer (RNN-T), and Token-and-Duration Transducer (TDT).
@@ 3.1 Connectionist Temporal Classification (CTC)
CTC (Graves et al., 2006) is a sequence-to-sequence prediction model designed to address the challenge of aligning input and output sequences of different lengths. Let $x=(x_{1},x_{2},...,x_{T})$ represent the input sequence, $y=(y_{1},y_{2},...,y_{U})$ the output sequence, and $\mathcal{V}$ the vocabulary, where $v\in\mathcal{V}$ includes all tokens in $y$ and a special “blank” symbol. A CTC model predicts $P(v|x_{t})$ for each time step $t$. The “blank” symbol is crucial for handling the discrepancy between input and output sequence lengths, allowing for many-to-one alignments and representing frames that do not contribute additional information to the output. CTC models employ a set of rules to generate CTC-augmented sequences that are equivalent to the original sequence: (1) When a token is repeated in the original sequence, at least one “blank” token must be inserted between the repeated tokens, and (2) Any token can be repeated an arbitrary number of times.
@@
CTC models with non-autoregressive components like self-attention and convolution are widely accepted as non-autoregressive models (Higuchi et al., 2021; 2020; Xu et al., 2023a), although during inference, CTC’s token deduplication and blank handling would still depend on the last emitted token. We point out that this NAR categorization is based on the neural network computation of the model. In CTC, all neural network computation during inference can be carried out non-autoregressively, and the only autoregression is done on basic data types (strings, lists, etc.), which induces negligible computation compared to the costly neural network computations. In this work, we also follow this principle in categorizing models as autoregressive or not.
@@ 3.2 Recurrent Neural Network Transducer (RNN-T)
The RNN-Transducer, proposed by Graves (2012), represents a significant advancement in sequence-to-sequence modeling. RNN-T comprises three key components: an encoder, a predictor, and a joint network. The encoder extracts high-level features from the audio input, while the predictor processes the text history. The joint network then combines these outputs to generate a probability distribution over the vocabulary.
@@
Despite its superior performance, RNN-T’s reliance on the predictor and text history introduces autoregressive behavior in inference. Consequently, RNN-T inference resists full parallelization, resulting in lower efficiency compared to CTC. This trade-off between performance and efficiency highlights the ongoing challenge of designing optimal sequence-to-sequence models for ASR.
@@ 3.3 Token-and-Duration Transducers (TDT)
Token-and-Duration Transducer (TDT) model is introduced by Xu et al. (2023b). Unlike traditional Transducers where only blank predictions increment the encoder index $t$ by exactly one, TDT supports advancing $t$ by multiple steps for both blank and non-blank predictions. A TDT model generates two probability distributions, one for token, and the other for durations, which determines the number of audio frames to skip. By supporting frame skipping, TDT models require significantly fewer decoding steps compared to traditional Transducers, resulting in faster inference.
@@ 4 Hybrid-Autoregressive Inference Transducer (HAINAN)
We propose Hybrid-Autoregressive Inference Transducers (HAINAN), an extension of the TDT model designed to support multiple modes of autoregression in inference. HAINAN incorporates stochastic predictor masking: during training, the predictor’s output is randomly masked before being passed to the joint network. We implement this with a masking probability of 0.5, sampled at the [batch, text-index] level. Except for the random masking performed in training, the rest of the model training of HAINAN (computation in other components, and loss computation, etc) remains identical to that of TDT training. As a result of this change, during training, half the time the joint network is learning to predict with the whole model, and the other half the time it is learning to predict with information from the encoder only. Therefore when the model is well-trained, the joint network can generate valid distributions regardless of whether the predictor output is provided.
The HAINAN model supports three inference modes: autoregressive (AR), non-autoregressive (NAR) depending on whether we use the predictor network during inference. In addition, we also propose semi-autoregressive (SAR) inference, which combines elements of both NAR and AR inference to achieve better speed-accuracy tradeoff, and a Viterbi-based inference, which leverages the better representations learned by HAINAN models for better accuracy.
Algorithm 1 Autoregressive Inference
1:  Input: acoustic input $x$
2:  enc = encoder($x$) # [T, H]
3:  hyp = []; $t$ = 0
4:  while $t<$ len(enc) do
5:     pred = predictor(hyp) # [H]
6:     token_probs, dur_probs = joint(enc[$t$], pred) # [V] and [D]
7:     token = argmax(token_probs)
8:     duration = argmax(dur_probs)
9:     if token is not blank then
10:        hyp.append(token)
11:     $t$ += duration
12:  Return hyp
Algorithm 2 Non-AR Inference
1:  Input: acoustic input $x$
2:  enc = encoder($x$) # [T, H]
3:  token_probs, dur_probs = parallel_joint(enc, pred=None) # [T, V] and [T, D]
4:  tokens = argmax(token_probs, dim=-1) # [T]
5:  durations = argmax(dur_probs,dim=-1) # [T]
6:  hyp = []; $t$ = 0
7:  while $t<$ len(enc) do
8:     token = tokens[t]
9:     if token is not blank then
10:        hyp.append(token)
11:     $t$ += max(1, duration[t]) # avoid infinite loop
12:  return hyp
4.1 Autoregressive inference
Autoregressive inference of HAINAN models is identical to TDT. For ease of reference, we provide Algorithm 1, which is a rewrite of the original Algorithm 2 from Xu et al. (2023b) with slightly changed notation/naming conventions to be consistent with this paper. To facilitate understanding of the Algorithms, we include tensor shapes as comments in the Algorithm, and we use T, V, D, U, H to represent audio-length, num-tokens, num-durations, text-index, and hidden-dimensions.
4.2 Non-autoregressive inference
Algorithm 2 shows NAR inference procedure. Compared to AR inference, NAR inference requires simple modifications of (1) passing an all-zero tensor instead of predictor output to the joint network; (2) moving the joint network computation out of the decoding loop, allowing parallel processing of all encoder outputs; (3) after each step, we increment $t$ by at least one to avoid infinite loops.
@@ 4.3 Semi-autoregressive inference
SAR inference is a novel algorithm combining elements of both AR and NAR approaches, as detailed in Algorithm 3. For ease of description, we assume the NAR inference procedure also returns the corresponding time-stamps of token outputs. The Algorithm consists of two main phases:
@@
    Hypothesis Refinement (lines 5-8) uses the predictor to compute representations of all partial history contexts in the current hypothesis. It then combines the predictor’s output with selected encoder output frames in the joint network to re-compute probability distributions for those frames, in parallel [^2] The computation at line 7 is done fully parallelized, and line 6 may require some sequential computation if the predictor is an LSTM. If the predictor is stateless, this computation can also be fully parallelized as well.. This procedure can be repeated to further improve the results. Note if repeated, all but the last round of the refinement should limit the argmax on line 9 to non-blank tokens since the token would need to be passed in the predictor for the next round of refinement.
The SAR approach combines the strengths of AR and NAR methods, allowing high degrees of parallelization while leveraging text histories in predictor computation to improve model accuracy.
Algorithm 3 Semi-Autoregressive Inference of HAINAN Models
1:  Input: acoustic input $x$
2:  enc = encoder(x) # [T, H]
3:  hyp, time_stamps = NARInference(enc) # [U] and [U]
4:  useful_frames = enc[time_stamps,:] # [U, H]
5:  shifted_hyp = [BOS] + hyp[:-1] # [U]
6:  pred = predictor(shifted_hyp) # [U, H]
7:  token_probs, duration_probs = parallel_joint(useful_frames, pred) # of shapes: [U, V] and [U, D]
8:  hyp = argmax(token_probs, dim=-1) # [U]
9:  Return: hyp
4.4 Viterbi-decoding
In addition to AR/NAR and SAR inference, we also propose a Viterbi-based inference method for HAINAN models. Given an audio sequence of length $T$, we view the output of the NAR HAINAN model as a directed acyclic graph (DAG) with $T+1$ nodes, while the first $T$ correspond to frames and the last acts as a special end-of-sentence (eos) token. For each frame, we only consider the arg-max token probability, which represents the weight of the first $T$ nodes (the eos node has a weight 1). The duration outputs at each frame correspond to transition weights into neighboring nodes. For example, in our models with max-duration 8, each node has at most 8 incoming and 8 outgoing connections connecting to its neighbors.
With this setting, speech recognition on the input audio is equivalent to finding the best path of the DAG from the first node to the last, which can be efficiently solved by a simplified Viterbi decoding algorithm. Notably, Viterbi and SAR techniques are orthogonal, and after Viterbi decoding generates the tokens and time stamps, we can use SAR techniques to further improve the results. The Algorithm for Viterbi decoding is shown in Algorithm 4. For simplicity, this Algorithm returns a “backtrack” structure that encodes the selected frames found by the best-path algorithm. The Algorithm’s key is between lines 7 and 14, where nested loops run over (time, duration) pairs, and for each node, it finds the probability of the best path from the first node to that node using a dynamic programming algorithm. The Algorithm can be carried out efficiently with $O(TD)$ time complexity, where $T$ is the input sequence length and $D$ is the number of durations. In terms of space complexity, this Viterbi algorithm requires $O(T)$ (to store best-probs and backtrace pointers for each time index). Because duration 0 is not allowed in NAR modes, we assume there’s no duration 0 in $N$.
Algorithm 4 Viterbi Decoding of HAINAN Models
1:  Input: acoustic input $x$, supported durations $N$
2:  enc = encoder(x)
3:  token_probs, duration_probs = parallel_joint(enc, pred=None) # of shapes [T, V] and [T, D]
4:  best_prob_per_frame = max(token_probs, dim=-1) # of shape [T]
5:  best_prob = [0.0 for t in range(len(x) + 1)]
7:  for target in range(1, len(x) + 1) do
8:     for idx, n in enumerate(N) do
10:        alpha = best-prob[source]
11:        trans_prob = best_prob_per_frame[target] * duration_prob[source,idx]
12:        if alpha * trans_prob $>$ best_prob[target] then
13:           best_prob[target] = alpha * trans_prob
14:           backtrack[target] = source
15:  Return: backtrack